# LOLA: LLM-Assisted Online Learning Algorithm for Content Experiments

### v2

@@ 3.1.3 Performance Analysis of Prompt-based Approaches
Before presenting the results, we make note of two points. First, we do not find any significant difference between GPT-4 and GPT-4o; therefore, we only report the results from GPT-3.5 and GPT-4 for simplicity. Second, the GPT models can produce different answers even when given the exact same prompt. To mitigate this variability, we randomly select 1,000 significant pairs of headlines and conduct all our experiments on this same test set. Similarly, all the demonstrations are also randomly sampled and fixed for all experiments to reduce the variance of prompt outputs.
@@
The results of all the prompt-based experiments are shown in Table 5. A more detailed pairwise comparison and t-test analysis of all the different experiments are available in Web Appendix §$\S$§B.1. The second column of Table 5, $n_{demo}$, represents the number of demonstrations used for in-context learning, with $n_{demo}=0$ indicating a zero-shot prompting. Additionally, for all the in-context learning experiments, we consider two scenarios—one where the demonstrations are shown with ground truth labels and another where they are shown with flipped outcome labels. The third column, “Flipped Label” indicates whether the labels were flipped or not.
GPT Model $n_{demo}$ Flipped Label Accuracy [0.025 0.975]
GPT-3.5 0 0 54.20% 51.10% 57.10%
GPT-3.5 2 0 48.70% 45.50% 51.50%
GPT-3.5 2 1 50.85% 47.60% 53.70%
GPT-3.5 5 0 51.30% 48.20% 54.30%
GPT-3.5 5 1 50.10% 46.80% 53.20%
GPT-4 0 0 57.26% 54.20% 60.00%
GPT-4 2 0 63.60% 60.50% 66.40%
GPT-4 2 1 62.90% 59.70% 65.70%
GPT-4 5 0 60.80% 57.70% 63.40%
GPT-4 5 1 61.80% 58.50% 64.80%
Table 5: Results for prompt-based approaches with different GPT models, the number of demonstrations, and whether the demo’s label is flipped or not. Accuracy is defined as the ratio of correct answers divided by the number of prompts (1000 in our analysis). The 95% confidence interval is calculated by performing 5,000 bootstrap resamples of the entire 1,000 pairs. To save on the monetary cost of prompts and not introduce additional variance, we did not rerun the GPT prompts for the same samples; instead, we used the responses/predictions from the original run.
@@
We find that prompt-based approaches using GPT-3.5, except for the zero-shot prompting, are not significantly better than random predictions (accuracy rates hover around 50%). Neither the number of demonstrations used nor the label type have any significant impact on the accuracy of the GPT-3.5 predictions. Interestingly, zero-shot prompting using GPT-3.5 is slightly better than random prediction and also marginally outperforms in-context learning. In comparison, experiments with GPT-4 have accuracy rates around 60-65%, which is a significant improvement in performance compared to GPT-3.5. For GPT-4, in-context learning significantly improves performance over zero-shot prompting in GPT-4, though increasing the number of demonstrations from two to five does not result in a significant difference in performance. There is also no significant difference between using correct labels and incorrect labels in the demonstrations for in-context learning. This confirms that in-context learning helps GPT understand the task better through the input-output examples, but it fails to leverage the information in mapping from inputs to outputs effectively.
@@
In summary, we find that prompt-based methods have low accuracy (no more than 65%), high randomness (different answer outputs even under the same prompt inputs), slow running time (compared to other methods introduced later in this section), and high monetary cost (see Web Appendix §$\S$§E for a detailed discussion of costs and running time). All of these drawbacks make them unsuitable for real-time business applications, i.e., these methods cannot replace experimentation for content selection in digital platforms.
3.2 Classification Models with LLM-based Text Embeddings
@@
We now consider another LLM-based method for the prediction task. Specifically, we leverage the text embeddings from OpenAI and utilize classification models, including Logistic Regression (Logit) and Multilayer Perceptron (MLP), to predict which of the headlines in a pair is likely to generate a higher CTR.
@@
Text embedding techniques transform large chunks of text, such as sentences, paragraphs, or documents, into numerical vectors. These embedding vectors can then be used in a variety of applications, including text classification, where texts are grouped into categories; semantic search, which improves search accuracy by understanding the meaning behind the search queries; and sentiment analysis, which is used to determine the emotional tone of a piece of text. [^7] Embedding models and standard prompt-based LLMs used share some common features, e.g., the underlying Transformer architectures and extensive pretraining on large text corpora. However, they are optimized for different tasks. Models like GPT are optimized for generating coherent and contextually relevant text based on input prompts while embedding models focus on producing high-quality embeddings for downstream tasks. See patil2023survey for a more detailed survey of the most recent text embedding models and their applications.
@@
An interesting question is why text embeddings obtained from LLM models trained for the next word/token prediction task with the cross-entropy loss are able to produce informative features for downstream classification tasks. (saunshi2020mathematical) provide an explanation by demonstrating that classification tasks can be reformulated as sentence completion tasks, which can be further solved linearly using the conditional distribution over words following an input text. Thus, the embeddings, which can be roughly viewed as the next token probability generated from a next-word prediction task, inherently capture the contextual information and relationships between words, making them useful for various downstream tasks such as sentiment analysis and headline selection (as our experiments illustrate).
3.2.1 Implementation Details
Figure 3 illustrates the conceptual framework of our approach here. We concatenate the embeddings of two headline embeddings888 Because embeddings approximate the next token probability, one can also use the difference (subtraction) of two embedding vectors instead of the concatenation as the input to the downstream classification model. In our task, these two input types have no significantly different impact on the final classification accuracy. and apply the downstream classification models to predict the probability that the first headline will generate a higher CTR. We consider three embedding models:
@@ 3.2.1 Implementation Details
    OpenAI embeddings: At the time of this study, the latest and best embedding model available from OpenAI was text-embedding-3-large. These embeddings are obtained using the OpenAI embedding API in Python. This embedding model generates embeddings with up to 3,072 dimensions and supports a maximum token limit of 8,191, with its knowledge base extending up to September 2021. A notable feature of this model is its flexibility – it allows users to tailor the embedding dimension to their specific needs. [^9] Other models capable of producing text embeddings, like BERT (devlin2018bert) and Sentence-BERT (reimers2019sentence), have fixed embedding vector dimensions that remain constant once the model is trained. These dimensions cannot be changed or resized for different applications. However, OpenAI’s embedding model uses a technique that, while producing outputs of fixed length, encodes information at different granularities (kusupati2022matryoshka). This allows a single embedding to adapt to the computational constraints of downstream tasks. More specifically, downstream tasks can use only the first several dimensions of the original vector, and the information retained decreases smoothly as a shorter part of the vector is used. In our experiments, we consider both 3,072- and 256-dimensional embeddings to examine whether the size of the embedding vector size affects the performance of the downstream classification models. [^10] A higher dimension does not necessarily mean better performance. While a higher dimension can retain more information for each sample, it also means that the samples are spread more sparsely in a higher-dimensional space given a limited number of training samples. This sparsity can increase the possibility of over-fitting.
@@
    Llama-3-8b: Meta’s Llama-3 (metallama3) is regarded as one of the most advanced open-source LLMs to date (llama3modelcard). We obtain embeddings from Llama-3 with 8 billion parameters (Llama-3-8b). We fine-tune this exact model using LoRA in §$\S$3.3; therefore, using the embeddings from this model allows for a direct comparison of fine-tuning vs. embedding-based classification methods (see §$\S$3.4 for comparisons). Llama-3-8b does not produce embeddings explicitly designed to represent the semantic meaning of entire input sequences. To address this issue, we apply the technique described in (behnamghader2024llm2vec), which transforms the Llama-3-8b model into a more suitable text embedding model, outputting embeddings with a dimension of 4,096.
@@
Next, to assess the impact of model complexity on performance, we consider two classification models:
  • $\bullet$
    Logit model with linear features: This is one of the simplest binary classification models and predicts the probability that a given input belongs to a particular class by applying the logistic function to a linear combination of the input variables.
  • $\bullet$
    Multilayer perceptron: MLP is a powerful neural network architecture consisting of multiple linear and non-linear hidden layers that allow the model to learn complex patterns in the data. It employs a fully connected neural network with ReLU activation for hidden layers and Sigmoid transformation for the output probability, with dropout layers added to prevent over-fitting. We refer interested readers to Chapter 5 in (zhang2021dive) for more details. We train the MLP using the stochastic gradient descent algorithm to minimize binary cross-entropy and perform hyper-parameter tuning with Bayesian Optimization to determine the optimal configuration. Please see Web Appendix §$\S$§C.1 for additional details of the fine-tuning process and the architectures used in the final implementation.
Figure 3: The pipeline of the classification with the LLM text embeddings. Headline 1 and headline 2 are natural language sentences while Embedding 1 and Embedding 2 are numerical vectors.
Finally, in terms of the data, as discussed earlier, we choose all headline pairs that are significant at the 0.05 level. Note that the same headline can appear in multiple pairs when there are more than two headlines tested in one experiment. To avoid information leakage, we first randomly split the data into 80% training and 20% test sets based on the test ID (the experiment ID for Upworthy) to ensure all headline pairs under the same experiment are either in the training set or the test set. Notice that Upworthy allowed additional experiments for the same headline under a new experiment ID because the editor may require an extra test to make an informed decision. It means we may have the same headlines in both the training and test sets, even if they were in different experiments. In such cases, we discard headlines that have appeared in the training set from the test set to avoid information leakage. [^11] There are alternative ways to do the train-test split without discarding any samples. One approach is assigning one headline pair at one time to a training or test set and then adding all pairs with the same headline in the same group until all samples are allocated. However, after conducting several experiments in our study, we found that the splitting procedure has minimal impact on the final results. Hence, to maintain clarity and simplicity in our exposition, we simply discard around 25% of test headlines that have appeared in the training set.
3.2.2 Performance Analysis of Embedding-based Approaches
In Table 6, we present the performance of various embedding-based classification models.
Embedding Model Classification Model Accuracy
OpenAI-256E Logit 77.31%
MLP 77.62%
OpenAI-3072E Logit 82.50%
MLP 82.26%
Llama-3-8b-4096E Logit 74.03%
MLP 79.40%
Word2Vec-256E Logit 67.89%
MLP 67.32%
Word2Vec-3072E Logit 67.55%
MLP 67.02%
Table 6: The performance of different embedding and classification models on test data. All models use the same training data, and accuracy is evaluated on the test data. Note that the Logit model’s training is a deterministic process after fixing the data, while MLP training incurs randomness via stochastic gradient descents. Therefore, we rerun MLP training three times with different sample paths (training set shuffling) and report the average accuracy from three reruns.
At a high level, we see that all the embedding-based classification models perform better than prompt-based approaches. This performance can be attributed to the fact that these models are trained on a customized dataset, allowing them to learn decision rules tailored to the headline selection task. Because these models can leverage a larger labeled dataset during training, they are more generalizable and robust. In contrast, in-context learning relies on the LLM’s ability to understand and generalize from only a few demonstrations. Furthermore, the task of choosing the catchier headline from two options may not be a common scenario in OpenAI’s training corpus. Consequently, GPT models may struggle to generalize effectively from the limited examples provided in prompts, leading to less accurate predictions.
Next, we compare performance under different embedding models. We see that OpenAI embeddings outperform both Llama-3-8b and Word2Vec-based models. This is understandable since OpenAI’s LLMs are state-of-the-art models that naturally perform better than the classic Word2Vec embeddings. Also, the performance of embeddings from open-source Llama-3-8b falls between that of the proprietary OpenAI model and the older Word2Vec that does not use the transformer architecture. We also see that for OpenAI, higher-dimensional embeddings give higher-quality encoding and, thus, higher prediction accuracy. However, there is some limit beyond which there are diminishing returns. In our experiment, the 3072-dimensional embedding (OpenAI-3072E) improves over the 256-dimensional embedding (OpenAI-256E) by around 5%. A more striking observation is that classification model complexity does not improve the prediction for OpenAI embeddings. Note that there is no significant difference in accuracy between MLP and Logit for the following embedding models – OpenAI-256E, OpenAI-3072E, Word2Vec-256E, and Word2Vec-3072E. But for Llama-3-8b embeddings, the MLP classification model outperforms Logit classification model by approximately 5%. We hypothesize that with more dimensions in text embeddings, the separation between different classes becomes more complex (e.g., higher sparsity and higher risk of over-fitting), requiring a more sophisticated model for classification. Overall, our findings suggest that a Logit classifier with linear features on top of fixed text embeddings from OpenAI can effectively capture the attractiveness of content.
We now investigate a few other aspects of these models and summarize our findings below. For brevity, we simply discuss the findings here and refer interested readers to the Web Appendix for details. Note that in all these analyses, we focus on OpenAI embeddings (since it offers the best performance) and the logistic regression with linear features as our classification model (since the MLP does not offer any significant improvement over the logit model for OpenAI embeddings and is easier to train).
  • $\bullet$
    First, we examine the conditions under which these approaches fail. We use cosine similarity measures to quantify the similarity between the two headlines in a pair and find that these approaches tend to perform worse when the headlines are similar. This is understandable since text embeddings may not be able to capture the differences between two headlines when they are similar. As a result, it is hard for a classification model to differentiate between them. See Web Appendix §$\S$§C.2 for details.
  • $\bullet$
    Second, we quantify the role of the number of training samples on predictive accuracy. For the 256E-dimensional embedding, we find that approximately 65% of the training set, comprising approximately 25,000 pairs (with the remaining 35% used for the test set), is adequate to achieve high and stable performance. However, when we use the 3072E-dimensional embedding, there is evidence of overfitting even when we use 95% of the data for training. This suggests that for larger embedding vectors, we need much larger training datasets. See Web Appendix §$\S$§C.3 for details.
  • $\bullet$
    Third, we examine whether we can improve the performance and alleviate any overfitting issues by adopting dimensionality reduction techniques (PCA) and/or regularization (adding $L^{1}$/$L^{2}$ norms). However, we don’t see any significant improvement with these methods. See Web Appendix §$\S$§C.4 for details.
  • $\bullet$
@@ 3.2.2 Performance Analysis of Embedding-based Approaches
    Finally, there are two potential threats to the inference from our analysis. First, one may be concerned as to whether Upworthy’s data has been used as part of OpenAI’s training corpus, which could potentially invalidate the results. A second concern is that readers’ preferences changed over time and that the current sample splitting method fails to take into account when the headline was created. We examine both these issues and provide a detailed discussion and robustness checks to alleviate concerns around them in Web Appendix §$\S$§C.5.
@@ 3.3 Fine-Tuning Open-Source LLMs with LoRA
In §$\S$3.2, we trained classification models using text embeddings as fixed inputs, bypassing the need to train the LLM itself. The best-performing embedding-based model reaches an accuracy of 82.50% without modifying any parameters in the LLM. This raises a natural question: if we can train LLMs directly, can we achieve better performance? However, training LLMs from scratch is costly and resource-intensive. Therefore, in this section, we turn to fine-tuning techniques, which involve partially adjusting the parameters in the LLM to improve performance while keeping the computational cost manageable. Note that to fine-tune LLMs, we need to be able to access the underlying model parameters. This is only feasible for open-source LLMs, like BERT and Llama. [^12] Some proprietary LLMs offer black-box fine-tuning. For instance, OpenAI allows fine-tuning GPT-3.5, GPT-4/GPT-4o (with limited experimental access) through its API in a black-box manner. Users provide the training and test datasets and select the base GPT model, but are not provided any details of the fine-tuning process or the underlying model parameter. This approach also has a number of other drawbacks – the fine-tuning API is quite expensive compared to the open-source fine-tuning approaches, the fine-tuned model cannot be hosted on a local machine, data privacy concerns since the media firm’s proprietary data will need to be submitted to OpenAI. Given these drawbacks, it is preferable for firms to fine-tune open-source LLMs, as the process and data are fully under the control of the firm. However, for the completeness of comparison, we fine-tuned the ChatGPT-3.5 model on the training set using this API, with the default configuration provided by OpenAI, and the same training-test dataset split. We obtained an accuracy of only 79.80% on the test set. The goal of fine-tuning is to adapt a general-purpose LLM for specific downstream tasks through a process typically involving supervised learning using labeled datasets.
@@
For the fine-tuning task, we use Meta’s Llama-3-8b as our primary base model. zhao_etal_2024 consider a large-scale experiment where they fine-tune different open-source LLMs and show that for NLP tasks, including news headline generation, Llama-3 offers one of the best performance. Llama-3 offers two versions: one with 8 billion parameters (Llama-3-8b) and another with 70 billion parameters (Llama-3-70b). We opted for the smaller version since it is easier to fine-tune and, as we will see, even with this smaller model, we see evidence of over-fitting (see §$\S$3.3.3 for details). Hence, we avoid fine-tuning the 70-billion-parameter model, where over-fitting problems are likely to be exacerbated.
The rest of this section is organized as follows. We present the LoRA method for efficient fine-tuning in §$\S$3.3.1. The goal is not to delve into every technical detail but to provide the necessary background information for a high-level understanding of fine-tuning. Then we detail our implementation in §$\S$3.3.2 and report our results in §$\S$3.3.3. Background information related to Transformer block, the fundamental component of most LLMs, including Llama-3, and the basic architecture of Llama-3 are given in Web Appendix §$\S$§D.1.
3.3.1 LoRA Fine-Tuning Technique
@@
Traditional fine-tuning of LLMs involves updating a substantial number of parameters, which can be computationally expensive and memory-intensive. To efficiently fine-tune the Llama-3 model on hardware with limited GPU memory, we employ Low-Rank Adaptation (LoRA), a popular Parameter-Efficient Fine-Tuning (PEFT) method developed by (hu2021lora), enabling practical fine-tuning of large-scale LLMs on medium-scale hardware. [^13] There are many alternatives for PEFT, like Prefix Tuning (li2021prefix), Adapter Layers (houlsby2019parameter), and BitFit (zaken2021bitfit). We select LoRA due to its excellent balance between simplicity and efficiency, making it the top choice for fine-tuning. LoRA operates under the assumption that updates during model adaptation exhibit an intrinsic low-rank property and introduce a set of low-rank trainable matrices into each layer of the Transformer model. This approach leverages the low-rank decomposition, which significantly reduces both memory usage and computational time through a significantly reduced number of trainable parameters.
We illustrate LoRA in Figure 4. We use $W$ to represent any pre-trained weight matrix, which can be either $W_{Q}$, $W_{V}$, or $W_{K}$. Specifically, $W_{Q}$ is the weight matrix for the query vectors, $W_{K}$ is for the key vectors, and $W_{V}$ is for the value vectors in the attention mechanism of the Transformer architecture. These matrices store most of the information and knowledge learned by LLM, and are crucial for calculating the attention scores and subsequently determining the relevance of different tokens in the input sequence. For more technical background, we refer to Web Appendix §$\S$§D.1. We use $d\times d_{W}$ to represent the dimension of the weight matrix and $r$ to represent the rank of the low-rank matrices. We denote the modified weight matrices after fine-tuning as $W^{\prime}$, which can be further expressed as:
Figure 4: Illustration of LoRA Fine-Tuning.
3.3.2 Implementation Details
We use exactly the same training and test sets as in §$\S$3.2, but the inputs here are the original headlines with the text information. To help Llama-3-8b understand the boundary between two headlines, we concatenate them using the default end of sequence (EOS) token, <|begin_of_text|>, placed between them. For example, for the headline pair The First Headline and The Second Headline, the combined sequence would be, <|begin_of_text|>The First Headline<|begin_of_text|>The Second Headline.
We add a classification head to the original Llama-3-8b base model, which includes 32 Transformer layers, as depicted in Figure 5. Llama-3-8b employs a causal attention mask, where the attention is unidirectional—each token can only attend to preceding tokens in the sequence. Thus, the last token in the sequence, being informed by all prior tokens, is used for classification. This embedding vector of the last token (which is 4096-dimensional), output by the last Transformer layer, is processed through a linear layer producing a two-dimensional logit vector, denoted by $Z=(Z_{1},Z_{2})$. The $Z$ vector is then transformed into probabilities by the softmax function, indicating the likelihood of each headline being catchier, i.e.,
$\mathbb{P}\{\text{i-th headline is catchier}\}=\frac{e^{Z_{i}}}{e^{Z_{1}}+e^{Z_{2}}}\mathrm{~{}for~{}}i=1,2.$
Figure 5: The classification pipeline using the Llama model. Headlines are first tokenized and each token is converted into an embedding vector. For illustrative purposes, only two Transformer layers are shown, although the actual model comprises 32 layers for the 8 billion parameter configuration. The figure highlights the use of uni-directional attention in Llama, whereby each token can only incorporate information from preceding tokens. Connections between Transformer blocks represent the attention mechanism, and only a subset of these links are displayed for simplicity. The output from the last token is fed into a classification head, which computes the probability that one headline is more impactful than the other.
During fine-tuning, our objective is to minimize the cross-entropy loss on the training set. We implement LoRA fine-tuning using the PEFT library developed by Hugging Face. We set the dimension of the low-rank matrices as $r=4$. Although LoRA can be applied to any subset of weight matrices in LLMs, applying LoRA to query weight matrices $W_{Q}$ and value weight matrices $W_{V}$ is a common practice and typically gives better performance (hu2021lora). Following this practice, in our experiment, we specifically apply LoRA to the $W_{Q}\in\mathbb{R}^{d\times d_{q}}$ and $W_{V}\in\mathbb{R}^{d\times d_{v}}$ matrices within the Transformer blocks of Llama-3-8b, where $d=4096,d_{q}=4096,d_{v}=1024$. To mitigate over-fitting issues, we also apply a relatively large dropout rate of 0.3 to these low-rank matrices. [^14] Furthermore, we enable 16-bit mixed precision training. Although 16-bit training slightly reduces numerical precision, it accelerates operations on GPUs, which enhances computational efficiency significantly. This mode makes the training process faster and consumes less memory compared to the default 32-bit training, especially for large-scale models like Llama-3-8b.

### v3

@@ 3.1.3 Performance Analysis of Prompt-based Approaches
Before presenting the results, we first make a note that we do not find any significant difference between GPT-4 and GPT-4o; therefore, we only report the results from GPT-3.5 and GPT-4 for simplicity. The results of all the prompt-based experiments are shown in Table 4. A more detailed pairwise comparison and t-test analysis of all the different prompts are available in Web Appendix §$\S$§B.1. The second column of Table 4, $n_{demo}$, represents the number of demonstrations used for in-context learning, with $n_{demo}=0$ indicating a zero-shot prompting. Additionally, for all the in-context learning experiments, we consider two scenarios—one where the demonstrations are shown with ground truth labels and another where they are shown with incorrect labels. The third column, “Incorrect Label” indicates whether the labels were correct (“Incorrect Label” = 0) or incorrect (“Incorrect Label” = 1).
Model $n_{demo}$ Incorrect Label Accuracy on the test set
Random Guess - - 33.02%
GPT-3.5 0 - 29.76%
GPT-3.5 2 0 34.23%
GPT-3.5 2 1 31.60%
GPT-3.5 5 0 30.95%
GPT-3.5 5 1 30.59%
GPT-4 0 - 37.85%
GPT-4 2 0 39.96%
GPT-4 2 1 38.77%
GPT-4 5 0 39.17%
GPT-4 5 1 38.28%
Table 4: Results for prompt-based approaches with different GPT models, the number of demonstrations, and whether we intentionally provide incorrect solutions in the demonstration or not. Accuracy is defined as the ratio of correct answers over the test set size. The best performance on all test data is boldfaced.
@@
We find that prompt-based approaches using GPT-3.5, are not significantly better than random guess or even slightly worse. In comparison, experiments with GPT-4 have accuracy rates around 38-40%, which is a significant improvement in performance compared to GPT-3.5, and all settings for GPT-4 outperform the random guess. For GPT-4, in-context learning with two demonstrations slightly improves performance over zero-shot prompting in GPT-4, though increasing the number of demonstrations from two to five does not result in a significant difference in performance. The difference between using correct labels and incorrect labels in the demonstrations is also not significant for in-context learning. This confirms that in-context learning helps GPT understand the task better through the input-output examples, but it fails to leverage the information in mapping from inputs to outputs effectively.
@@
In summary, we find that prompt-based methods have low accuracy (no more than 40%), slow running time (compared to other methods introduced later in this section), and high monetary cost (see Web Appendix §$\S$§E for a detailed discussion of costs and running time). All of these drawbacks make them unsuitable for real-time business applications, i.e., these methods cannot replace experimentation for content selection in digital platforms.
3.2 CTR Prediction Models with LLM-based Text Embeddings
@@
We now consider another LLM-based method for the headline selection task. Specifically, we leverage the text embeddings from OpenAI and utilize regression models, including Linear Regression (Linear) and Multilayer Perceptron (MLP), to predict the CTRs of headlines for the same article and then select the one with the highest predicted CTR as the winning headline.
@@
Text embedding techniques transform large chunks of text, such as sentences, paragraphs, or documents, into numerical vectors. These embedding vectors can then be used in a variety of applications, including text classification, where texts are grouped into categories; semantic search, which improves search accuracy by understanding the meaning behind the search queries; and sentiment analysis, which is used to determine the emotional tone of a piece of text. [^8] Embedding models and standard prompt-based LLMs share some common features, e.g., the underlying Transformer architectures and extensive pretraining on large text corpora. However, they are optimized for different tasks. Models like GPT are optimized for generating coherent and contextually relevant text based on input prompts while embedding models focus on producing high-quality embeddings for downstream tasks. See patil2023survey for a more detailed survey of the most recent text embedding models and their applications.
@@
An interesting question is why text embeddings obtained from LLM models trained for the next word/token prediction task with the cross-entropy loss are able to produce informative features for downstream tasks. (saunshi2020mathematical) provide an explanation by demonstrating that typical downstream tasks can be reformulated as sentence completion tasks, which can be further solved linearly using the conditional distribution over words following an input text. Thus, the embeddings, which can be roughly viewed as the next token probability generated from a next-word prediction task, inherently capture the contextual information and relationships between words, making them useful for various downstream tasks such as sentiment analysis and headline selection (as our experiments illustrate).
Figure 3 illustrates the conceptual framework of our approach. We first use the LLM-based embedding model to transform the original headlines into embeddings. Then, we use the embeddings as inputs and CTRs as outputs to train the downstream CTR prediction model. For the test set, we leverage the trained regression model to predict the CTR of one headline at a time. After getting all CTR predictions of headlines in an A/B test, we rank the headlines with the highest CTR as the winning headline. We consider four embedding models, including the OpenAI embedding model, Llama-3-8b, Word2Vec, and BERT. Note that we do not train the embedding model and only train the downstream CTR prediction model using the training dataset, containing 12,376 headline A/B tests. A detailed introduction to embedding models and the implementation specifics are provided in Web Appendix §$\S$§C.1.
Figure 3: The pipeline of the headline selection using LLM text embeddings. We use an A/B test with three headlines for illustration. Headlines 1, 2, and 3 are natural language sentences, while Embedding 1, 2, and 3 are numerical vectors.
3.2.1 Performance Analysis of Embedding-based Approaches
Embedding Model CTR Prediction Accuracy on the test data
OpenAI-256E Linear 46.28%
MLP 44.38%
OpenAI-3072E Linear 43.06%
MLP 45.17%
Table 5: The performance of OpenAI embeddings and CTR prediction models on the test data. All models use the same training data, and accuracy is evaluated on the test data. Note that the Linear model’s training is a deterministic process after fixing the data, while MLP training incurs randomness via stochastic gradient descents. Therefore, we rerun MLP training three times with different random seeds and report the average accuracy from three reruns.
In Table 5, we present the performance of CTR prediction models based on OpenAI’s embeddings, including the 256-dimensional embedding (OpenAI-256E) and the 3072-dimensional embedding (OpenAI-3072E, which is the maximum size). The full performance results, including Llama-3-8b, Word2Vec, and BERT, can be found in Web Appendix §$\S$§C.2. At a high level, we see that all the embedding-based regression models obtain better accuracy than prompt-based approaches. This performance can be attributed to the fact that these models are trained on a customized dataset, allowing them to learn decision rules tailored to the headline selection task. Because these models can leverage a larger labeled dataset during training, they are more generalizable and robust. In contrast, in-context learning relies on the LLM’s ability to understand and generalize from only a few demonstrations. Furthermore, the task of choosing the best headline from multiple candidates may not be a common scenario in OpenAI’s training corpus. Consequently, GPT models may struggle to generalize effectively from the limited examples provided in prompts, leading to less accurate predictions.
We also find that, for OpenAI, higher-dimensional embeddings do not lead to higher prediction accuracy. Given that the 256-dimensional embedding stores only essential information, while the 3072-dimensional embedding also captures finer details, this performance gap is possibly due to the overfitting with the limited size of the training set. Predictions using higher-dimensional embeddings may cause the prediction model to focus too much on details and overfit the training set. A more striking observation is that increasing the complexity of the CTR prediction model does not necessarily improve predictions for OpenAI embeddings. We hypothesize that with more dimensions in text embeddings, the CTR prediction task requires a more sophisticated model due to the higher risk of overfitting. Overall, our findings suggest that a Linear CTR prediction model on top of fixed text embeddings from OpenAI can effectively capture the attractiveness of content.
We now investigate three other aspects of these models and summarize our findings below. For brevity, we simply discuss the findings here and refer interested readers to the Web Appendix for details.
@@ 3.2.1 Performance Analysis of Embedding-based Approaches
    We compare performance under different embedding models. We observe that OpenAI embeddings outperform Llama-3-8b, Word2Vec, and BERT-based models. See Web Appendix §$\S$§C.2 for details.
@@
    We quantify the impact of training sample size on predictive accuracy. In Web Appendix §$\S$§C.3, we train the model using varying proportions of the training data and evaluate its performance. We find that larger training datasets lead to higher accuracy rates, although the improvement is gradual. For instance, increasing the training size from 20% to 70% of the total dataset raises accuracy from 44% to 46% on the test data.
@@
    There are two potential threats to the inference from our analysis. First, one may be concerned as to whether Upworthy’s data has been used as part of OpenAI’s training corpus, which could potentially invalidate the results. A second concern is that readers’ preferences changed over time and that the current sample splitting method fails to take into account when the headline was created. We examine both these issues and provide a detailed discussion and robustness checks to alleviate concerns around them in Web Appendix §$\S$§C.4.
@@ 3.3 Fine-Tuning Open-Source LLMs with LoRA
In §$\S$3.2, we trained prediction models using text embeddings as fixed inputs, bypassing the need to train the LLM itself. The best-performing embedding-based model reaches an accuracy of 46.28% on the test set without modifying any parameters in the LLM. This raises a natural question: if we can train LLMs directly, can we achieve better performance? However, training LLMs from scratch is costly and resource-intensive. Therefore, in this section, we turn to fine-tuning techniques, which involve partially adjusting the parameters in the LLM to improve performance while keeping the computational cost manageable. Note that to fine-tune LLMs, we need to be able to access the underlying model parameters. This is only feasible for open-source LLMs like BERT and Llama. [^9] Some proprietary LLMs offer black-box fine-tuning. For instance, OpenAI allows fine-tuning GPT models through its API in a black-box manner. Users provide the training and test datasets and select the base GPT model, but are not provided any details of the fine-tuning process or the underlying model parameters. This approach also has a number of other drawbacks – the fine-tuning API is quite expensive compared to the open-source fine-tuning approaches, the fine-tuned model cannot be hosted on a local machine, data privacy concerns since the media firm’s proprietary data will need to be submitted to OpenAI. Given these drawbacks, it is preferable for firms to fine-tune open-source LLMs, as the process and data are fully under the control of the firm. However, for the completeness of comparison, we fine-tuned the GPT-4o (gpt-4o-2024-08-06 version) model on the training set using this API, with the default configuration provided by OpenAI and the same training and test datasets. Input to the model is several headlines of one article, and the target output is the index of the catchiest headline. We obtained an accuracy of 48.82% on the test set, but this result comes with a high cost of \$71.77 for the test set alone. The goal of fine-tuning is to adapt a general-purpose LLM for specific downstream tasks through a process typically involving supervised learning using labeled datasets.
@@
For the fine-tuning task, we use Meta’s Llama-3-8b as our primary base model. zhao_etal_2024 consider a large-scale experiment where they fine-tune different open-source LLMs and show that for NLP tasks, including news headline generation, Llama-3 offers one of the best performance. Llama-3 offers two versions: one with 8 billion parameters (Llama-3-8b) and another with 70 billion parameters (Llama-3-70b). We opted for the smaller version since it is easier to fine-tune and, as we will see, even with this smaller model, we see evidence of over-fitting (see §$\S$3.3.1 for details). Hence, we avoid fine-tuning the 70-billion-parameter model, where over-fitting problems are likely to be exacerbated. Background information related to the Transformer block, the fundamental component of most LLMs, including Llama-3, and the basic architecture of Llama-3 is given in Web Appendix §$\S$§D.1.
@@
Traditional fine-tuning of LLMs involves updating a substantial number of parameters, which can be computationally expensive and memory-intensive. To efficiently fine-tune the Llama-3 model on hardware with limited GPU memory, we employ Low-Rank Adaptation (LoRA), a popular Parameter-Efficient Fine-Tuning (PEFT) method developed by (hu2021lora), enabling practical fine-tuning of large-scale LLMs on medium-scale hardware. [^10] There are many alternatives for PEFT, like Prefix Tuning (li2021prefix), Adapter Layers (houlsby2019parameter), and BitFit (zaken2021bitfit). We select LoRA due to its excellent balance between simplicity and efficiency, making it the top choice for fine-tuning. LoRA operates under the assumption that updates during model adaptation exhibit an intrinsic low-rank property and introduce a set of low-rank trainable matrices into each layer of the Transformer model. This approach leverages the low-rank decomposition, which significantly reduces both memory usage and computational time through a significantly reduced number of trainable parameters. More details about LoRA fine-tuning and the implementation are given in Web Appendix §$\S$§D.2 and §$\S$§D.3 respectively.